# Aharonov-Bohm oscillations and perfectly transmitted mode in amorphous topological insulator nanowires

### v1

@@ I Introduction
A strong $3d$ topological insulator in the quasi-$1d$ geometry of a nanowire behaves differently than in a $3d$ geometry. In $3d$ the energy spectrum hosts an odd number of Dirac cones at the surface Zhang et al. (2009a) and is thus gapless. The electronic states at the surface are spin-momentum locked and protected from localization by the interplay of time-reversal symmetry and topology Bardarson et al. (2007); Nomura et al. (2007); Hasan and Kane (2010); Moore (2010); Qi and Zhang (2011); Bardarson and Moore (2013). In a topological nanowire the energy spectrum is gapped, and there is an even number of modes at every energy Rosenberg et al. (2010); Zhang et al. (2009b); Ostrovsky et al. (2010); Bardarson et al. (2010); Bardarson and Moore (2013); Bardarson and Ilan (2018). The gap is due to the spin of the electrons acquiring a $\pi$ Berry phase when taking a closed trajectory around the circumference of the nanowire Ostrovsky et al. (2010); Rosenberg et al. (2010); Bardarson et al. (2010); Zhang and Vishwanath (2010). Threading a magnetic flux $\phi$ through the cross section of the nanowire allows to change the number of modes from even to odd. At $\phi=n\phi_{0}/2$, with $\phi_{0}=e/h$ the flux quantum, the gap closes, restoring a Dirac cone and resulting in an odd number of modes at any Fermi energy. This occurs because the Aharanov-Bohm phase exactly cancels the spin Berry phase at these particular values of the magnetic flux Bardarson et al. (2010).
@@
The interplay between the number of modes and the time-reversal symmetry has consequences for the conductance of the nanowire Bardarson and Ilan (2018). Although the time-reversal symmetry is broken due to the magnetic flux, it is effectively recovered at the surface for $\phi=n\phi_{0}/2$. Intuitively, the symmetry results from electrons traveling in opposite directions around the circumference of the nanowire, picking up the same phase factor. The odd number of modes, together with time-reversal symmetry, implies that at $\phi=n\phi_{0}/2$ at least one reflection eigenvalue vanishes due to the antisymmetry of the reflection matrix Bardarson (2008). A perfectly transmitted mode protected from backscattering appears then in the nanowire. The conductance displays Aharanov-Bohm oscillations with varying magnetic flux and a quantized value at $\phi=n\phi_{0}/2$ Bardarson et al. (2010); Bardarson and Moore (2013); Zhang and Vishwanath (2010). The oscillations are a signature of the topological nature of the surface state in a $3d$ topological insulator Bardarson and Ilan (2018).
@@
Topological nanowires have been studied in experiments Peng et al. (2010); Xiu et al. (2011); Dufouleur et al. (2013); Hong et al. (2014); Cho et al. (2015); Jauregui et al. (2016); Dufouleur et al. (2017); Ziegler et al. (2018); Münning et al. (2021); Feng et al. (2025) and are potential candidates for technological applications, for example as electrical interconnects Gilbert (2021); Gupta et al. (2014); Philip et al. (2016); Rocchino et al. (2024); Khan et al. (2025); Gilbert (2025); Cheon et al. ; Soulié et al. (2024). The resistivity of traditional copper-based interconnects increases due to finite-size effects as their dimensions are reduced to the nanoscale Steinhögl et al. (2002); Steinlesberger et al. (2005). The absence of backscattering of topological surface states in nanowires may circumvent this problem. This strategy is already being investigated in topological-insulator Gupta et al. (2014); Philip et al. (2016) and Weyl-semimetal Rocchino et al. (2024); Khan et al. (2025); Cheon et al. ; Kaladzhyan and Bardarson (2019) nanowires. Another possible application is in the context of quantum computation. Analogously to semiconducting nanowires with strong spin-orbit coupling Oreg et al. (2010); Lutchyn et al. (2010); Alicea et al. (2011), topological nanowires in contact with a superconducting substrate can be brought to a topological superconducting state with a Majorana zero mode at their ends Kitaev (2001); Cook and Franz (2011); Cook et al. (2012); Juan et al. (2014); Manousakis et al. (2017); de Juan et al. (2019); Bai et al. (2022); Feng et al. (2025); Kaufhold and Rosch , which can be used as a topological qubit.
@@
The feasibility of these technological applications depends on the properties of topological nanowires being robust to, among other effects, disorder and structural imperfections that are common in real materials. Conversely, it is important to find materials that provide versatile and scalable platforms to create these nanowires. Amorphous topological materials Agarwala and Shenoy (2017); Xiao and Fan (2017); Mitchell et al. (2018); Pöyhönen et al. (2018); Mansha and Chong (2017); Costa et al. (2019); Agarwala et al. (2020); Marsal et al. (2020); Corbae et al. (2021); Muñoz Segovia et al. (2023); Grushin ; Corbae et al. (2023a) are one such platform, as they are easier to grow than perfectly crystalline structures. Recent experimental progress in which the surface state of amorphous $\mathrm{Bi}_{2}\mathrm{Se}_{3}$ Barzola-Quiquia et al. (2015); Banerjee et al. (2017) has been observed by ARPES measurements Corbae et al. (2023b); Ciocys et al. (2024) exemplifies their effectiveness as candidates for topological materials. When other topological materials are grown amorphous they still display transport effects often attributed to a non-zero Berry curvature DC et al. (2018); Ramaswamy et al. (2019); Bouma et al. (2020); Li et al. (2021); Molinari et al. (2023); Fujiwara et al. (2023); Zhao et al. (2025). Another example is the recent observation that electrical resistivity in thin films of noncrystalline CoSi and NbP is lower than in conventional metals Rocchino et al. (2024); Khan et al. (2025), potentially useful for nanoelectronic applications.
@@
Although promising, nanowires with an underlying amorphous structure can break the effective time-reversal symmetry of the surface due to the variation of their cross section Xypakis et al. (2020). It is thus pertinent to ask under which conditions the Aharanov-Bohm oscillations and perfectly transmitted mode survive in an amorphous nanowire. This question has been discussed for disordered nanowires Bardarson et al. (2010); Bardarson and Moore (2013); Xypakis and Bardarson (2017), but studies on amorphous nanowires remain scarce. Ref. Xypakis et al., 2020 studied nanowires with longitudinal ripples using a surface theory approach, Ref. Kozlovsky et al., 2020 studied cone-shaped wires, while Ref. Mal et al., focused on transport calculations in a tight-binding model of a diamond lattice with connectivity defects.
[…]
@@ II Model
We endow the amorphous nanowires with a Hamiltonian adapting Fu and Berg’s model Zhang et al. (2009a); Liu et al. (2010); Fu and Berg (2010) for a $3d$ topological insulator with spin-$1/2$ and two $p_{z}$ orbitals to the noncrystalline case. The resulting Hamiltonian reads
^$\begin{split}&\quad\qquad\hat{\mathcal{H}}=\sum_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\hat{c}^{\dagger}_{\boldsymbol{r}}h_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\hat{c}_{\boldsymbol{r}^{\prime}},\\ h_{\boldsymbol{r}\boldsymbol{r}^{\prime}}&=(\epsilon\tau_{1}-E_{F}\tau_{0})\delta_{\boldsymbol{r}\boldsymbol{r}^{\prime}}-t_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\tau_{1}-\dfrac{i}{2}\eta_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\tau_{2}\cos\theta\\ &+\dfrac{i}{2}\lambda_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\tau_{3}(\sigma_{2}\cos\varphi\sin\theta-\sigma_{1}\sin\varphi\sin\theta),\end{split}$ (1)
@@
and is in symmetry class AII Altland and Zirnbauer (1997); Kitaev (2009); Ludwig (2015) since $\mathcal{T}^{2}=-1$. In the crystalline limit, $\theta\in\{0,\pi/2,\pi\}$, $\phi\in\{0,\pm\pi/2,\pi\}$ and $r_{0}=1$, it hosts a strong topological phase for $\lambda,\eta>0$ and $2t<\epsilon<6t$. We use $t=1$, $\epsilon=4$, $\lambda=1$, $\eta=1.8$ and take $r_{0}=1.3$ in all numerical calculations.
@@
In order to compute the electronic transport we connect the nanowire to two semi-infinite translation-invariant leads described by the crystalline limit of Eq. (1). The leads are threaded by the same magnetic flux and have the same hopping and spin-orbit parameters as the nanowire, together with an additional chemical potential $\mu$ such that $E_{F}$ is replaced by $E_{F}-\mu$. 222 There is freedom to choose the value of the chemical potential. A nonzero chemical potential makes the density of states change abruptly from the leads to the nanowire, increasing interference and giving a highly oscillating conductance signal, see Fig. 2(a). We may then set $\mu=0$ to have a clearer and easier to interpret conductance signal, as in Fig. 1 and Fig. 6. In cases where the conductance is calculated at a fixed Fermi energy, the interference effects described above do not affect the results as much, and we may then use a nonzero chemical potential to allow the leads to have a higher density of states than the nanowire, acting as metallic reservoirs. This is the case in Fig. 2(d), Fig. 4(a) and Fig. 5. . The zero-bias conductance of the lead-nanowire-lead system is given by the Landauer formula Datta (1997)
@@
with $T$ the transmission matrix at the Fermi energy $E_{F}$. We numerically calculate $G$ with the software package Kwant Groth et al. (2014). To relate transport properties to the topology of the nanowires we use the local markers introduced in Ref. Hannukainen et al., 2022, which characterize the topology of states that break translation invariance and can be applied to amorphous materials Hannukainen et al. (2022, 2024).
The code and data used to perform the calculations discussed in this article is available in Ref. Martínez et al., 2025.
@@
Layered-amorphous nanowires are a useful starting point for understanding how amorphicity influences electronic transport in nanowires, as they maintain translation invariance along their axis. As a consequence, the magnetic flux is constant along the nanowire and the surface recovers an effective time-reversal symmetry at $\phi=n\phi_{0}/2$ that results in a perfectly transmitted mode.
@@
We demonstrate the presence of a perfectly transmitted mode by calculating the conductance of a single amorphicity realization as a function of the Fermi energy, shown in Fig. 1. In the absence of flux the conductance is zero until $E_{F}\simeq 0.06$, indicating the presence of an energy gap. At $\phi/\phi_{0}=0.56$ the conductance is quantized to $G=e^{2}/h$ for $E_{F}<0.12$, reflecting the perfectly transmitted mode. Note that the conductance is not exactly quantized down to $E_{F}=0$ because the leads have a finite size gap at $E_{F}=0$ when $\mu=0$. Since the surface state has a finite penetration length into the bulk, the flux it encloses deviates from the exact flux threaded through the full cross section. As a result, the perfectly transmitted mode does not occur exactly at $\phi/\phi_{0}=1/2$, but approaches this value in the limit of large $N_{x},N_{y}$  [^3] Apart from the shift induced by the finite bulk penetration depth, the position of the peaks may also be shifted to $\phi>n\phi_{0}/2$ if the amorphous cross section is smaller than reference cross section area $\mathcal{A}_{\rm ref}$, and to $\phi<n\phi_{0}/2$ if the amorphous cross section is bigger. This effect depends on the particular amorphicity realization.. Increasing the Fermi energy makes the flux value at which the conductance is maximal alternate between $\phi/\phi_{0}=0$ and $\phi/\phi_{0}=0.56$, as the number of modes at the Fermi energy changes between even and odd. The main difference from the crystalline case is the sharp conductance resonances seen especially at high Fermi energy. These come from the interface between the wire and the leads, where the lattice structure changes from being perfectly crystalline to being layered-amorphous, allowing electrons to backscatter and interfere. When the Fermi energy reaches the bands hosting states with a nonzero bulk contribution, scattering becomes stronger leading to an increase in interference and the upward trend in conductance gets flattened.
@@
In Fig. 2(a-c) we show the realization-averaged conductance of a layered-amorphous nanowire as a function of the Fermi energy and at three different amorphicities. As $w$ increases, the atomic arrangements of the different nanowire realizations become more distinct from each other. Since the exact details of each conductance realization depend on the particular nanowire structure, conductance curves dephase when averaged, leading to smoother curves than for small $w$. Close to $E_{F}=0$, the perfectly transmitted mode dominates the conductance.
@@
Figure 2: (a-c) Realization-averaged conductance of a layered-amorphous nanowire for three different amorphicities as a function of Fermi energy. The average is performed over $300$ layer amorphicity realizations of nanowires with $L=100$ and $N_{x}=N_{y}=10$. The chemical potential in the leads is $\mu=-1$. The shaded regions indicate the standard deviation of the distribution. (d) Single-realization conductance at $E_{F}=0$ as a function of magnetic flux and for different $w$. The parameters are $\mu=-1$, $L=200$ and $N_{x}=N_{y}=10$. (e) Band gap for the different nanowires in (d).
@@
Fig. 2(d) shows the single-realization conductance at $E_{F}=0$ as a function of magnetic flux, and Fig. 2(e) the associated band gap. Conductance peaks coincide with gap closings, deviating from the crystalline case as either $w$ or $\phi$ increase. Despite these deviations, the conductance peaks reach perfect transmission up to $w=0.2$. Note that this is not the case if we calculate the average conductance since the flux at which perfect transmission is attained changes between realizations, decreasing the height of the peak.

### v2

@@ I Introduction
A strong $3d$ topological insulator in the quasi-$1d$ geometry of a nanowire behaves differently than in a $3d$ geometry. In $3d$ the energy spectrum hosts an odd number of Dirac cones at the surface [1] and is thus gapless. The electronic states at the surface are spin-momentum locked and protected from localization by the interplay of time-reversal symmetry and topology [2; 3; 4; 5; 6; 7]. In a topological nanowire the energy spectrum is gapped, and there is an even number of modes  [^1] Throughout the manuscript we use the term mode to refer to electronic states with a particular sign of momentum. For example, if there exists one electronic state with energy $E$ and $k<0$ and another state with the same energy and $k>0$, there is only one mode as for each sign of momentum there is only one state. at every energy [9; 10; 11; 12; 7; 13]. The gap is due to the spin of the electrons acquiring a $\pi$ Berry phase when taking a closed trajectory around the circumference of the nanowire [11; 9; 12; 14]. Threading a magnetic flux $\phi$ through the cross section of the nanowire allows to change the number of modes from even to odd. At $\phi=\phi_{0}/2$ and up to integer multiples of the flux quantum $\phi_{0}=e/h$, the gap closes, restoring a Dirac cone and resulting in an odd number of modes at any Fermi energy. This occurs because the Aharanov-Bohm phase exactly cancels the spin Berry phase at these particular values of the magnetic flux [12].
@@
The interplay between the number of modes and the time-reversal symmetry has consequences for the conductance of the nanowire [13]. Although the time-reversal symmetry is broken due to the magnetic flux, it is effectively recovered at the surface for $\phi=(2n+1)\phi_{0}/2$. Intuitively, the symmetry results from electrons traveling in opposite directions around the circumference of the nanowire, picking up the same phase factor. The odd number of modes, together with time-reversal symmetry, implies that at $\phi=(2n+1)\phi_{0}/2$ at least one reflection eigenvalue vanishes due to the antisymmetry of the reflection matrix [15]. A perfectly transmitted mode protected from backscattering appears then in the nanowire. The conductance displays Aharanov-Bohm oscillations with varying magnetic flux and a quantized value at $\phi=(2n+1)\phi_{0}/2$ [12; 7; 14]. The oscillations are a signature of the topological nature of the surface state in a $3d$ topological insulator [13].
@@
Topological nanowires have been studied in experiments [16; 17; 18; 19; 20; 21; 22; 23; 24; 25] and are potential candidates for technological applications, for example as electrical interconnects [26; 27; 28; 29; 30; 31; 32; 33]. The resistivity of traditional copper-based interconnects increases due to finite-size effects as their dimensions are reduced to the nanoscale [34; 35]. The absence of backscattering of topological surface states in nanowires may circumvent this problem. This strategy is already being investigated in topological-insulator [27; 28] and Weyl-semimetal [29; 30; 32; 36] nanowires. Another possible application is in the context of quantum computation. Analogously to semiconducting nanowires with strong spin-orbit coupling [37; 38; 39], topological nanowires in contact with a superconducting substrate can be brought to a topological superconducting state with a Majorana zero mode at their ends [40; 41; 42; 43; 44; 45; 46; 25; 47], which can be used as a topological qubit.
@@
The feasibility of these technological applications depends on the properties of topological nanowires being robust to, among other effects, disorder and structural imperfections that are common in real materials. Conversely, it is important to find materials that provide versatile and scalable platforms to create these nanowires. Amorphous topological materials [48; 49; 50; 51; 52; 53; 54; 55; 56; 57; 58; 59] are one such platform, as they are easier to grow than perfectly crystalline structures. Recent experimental progress in which the surface state of amorphous $\mathrm{Bi}_{2}\mathrm{Se}_{3}$ [60; 61] has been observed by ARPES measurements [62; 63] exemplifies their effectiveness as candidates for topological materials. When other topological materials are grown amorphous they still display transport effects often attributed to a non-zero Berry curvature [64; 65; 66; 67; 68; 69; 70]. Another example is the recent observation that electrical resistivity in thin films of noncrystalline CoSi and NbP is lower than in conventional metals [29; 30], potentially useful for nanoelectronic applications.
@@
Although promising, nanowires with an underlying amorphous structure can break the effective time-reversal symmetry of the surface due to the variation of their cross section [71]. It is thus pertinent to ask under which conditions the Aharanov-Bohm oscillations and perfectly transmitted mode survive in an amorphous nanowire. This question has been discussed for disordered nanowires [12; 7; 72], but studies on amorphous nanowires remain scarce. Ref. 71 studied nanowires with longitudinal ripples using a surface theory approach, Ref. 73 studied cone-shaped wires, while Ref. 74 focused on transport calculations in a tight-binding model of a diamond lattice with connectivity defects.
In this work, we perform transport calculations on amorphous nanowires, both in the case of a constant cross section along the axis and in a fully amorphous case where there is no notion of a well-defined cross section. We use a tight-binding model with a crystalline limit describing the family of $\mathrm{Bi}_{2}\mathrm{Se}_{3}$ materials, one of the most prominent examples of 3dd topological insulators both in crystalline and amorphous forms. We investigate the presence of a perfectly transmitted mode and Aharanov-Bohm oscillations, and relate their robustness to the topological phase of the bulk of the nanowires by using local topological markers [75; 76; 77; 78; 79; 80; 81].
@@ II Model
We endow the amorphous nanowires with a Hamiltonian adapting Fu and Berg’s model [1; 82; 83] for a $3d$ topological insulator with spin-$1/2$ and two $p_{z}$ orbitals to the noncrystalline case. The resulting Hamiltonian reads
$\begin{split}&\quad\qquad\hat{\mathcal{H}}=\sum_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\hat{c}^{\dagger}_{\boldsymbol{r}}h_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\hat{c}_{\boldsymbol{r}^{\prime}},\\ h_{\boldsymbol{r}\boldsymbol{r}^{\prime}}&=(\epsilon\tau_{1}-E_{F}\tau_{0})\delta_{\boldsymbol{r}\boldsymbol{r}^{\prime}}-t_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\tau_{1}-\dfrac{i}{2}\eta_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\tau_{2}\cos\theta\\ &+\dfrac{i}{2}\lambda_{\boldsymbol{r}\boldsymbol{r}^{\prime}}\tau_{3}(\sigma_{2}\cos\varphi\sin\theta-\sigma_{1}\sin\varphi\sin\theta),\end{split}$ (1)
@@
and is in symmetry class AII [85; 86; 87] since $\mathcal{T}^{2}=-1$. In the crystalline limit, $\theta\in\{0,\pi/2,\pi\}$, $\phi\in\{0,\pm\pi/2,\pi\}$ and $r_{0}=1$, it hosts a strong topological phase for $\lambda,\eta>0$ and $2t<\epsilon<6t$. We use $t=1$, $\epsilon=4$, $\lambda=1$, $\eta=1.8$ and take $r_{0}=1.3$ in all numerical calculations unless explicitly stated.
@@
In order to compute the electronic transport we connect the nanowire to two semi-infinite translation-invariant leads described by the crystalline limit of Eq. (1). The leads are threaded by the same magnetic flux and have the same hopping and spin-orbit parameters as the nanowire, together with an additional chemical potential $\mu$ such that $E_{F}$ is replaced by $E_{F}-\mu$.  [^3] There is freedom to choose the value of the chemical potential. A nonzero chemical potential makes the density of states change abruptly from the leads to the nanowire, increasing interference and giving a highly oscillating conductance signal, see Fig. 2(a). We may then set $\mu=0$ to have a clearer and easier to interpret conductance signal, as in Fig. 1 and Fig. 6. In cases where the conductance is calculated at a fixed Fermi energy, the interference effects described above do not affect the results as much, and we may then use a nonzero chemical potential to allow the leads to have a higher density of states than the nanowire, acting as metallic reservoirs. This is the case in Fig. 2(d), Fig. 4(a) and Fig. 5. . The zero-bias conductance of the lead-nanowire-lead system is given by the Landauer formula [89]
@@
with $T$ the transmission matrix at the Fermi energy $E_{F}$, which relates the amplitudes of incoming and outgoing modes at the leads. We numerically calculate $G$ with the software package Kwant [90]. To relate transport properties to the topology of the nanowires we use the local markers introduced in Ref. 80, which characterize the topology of states that break translation invariance and can be applied to amorphous materials [80; 81].
@@
Layered-amorphous nanowires are a useful starting point for understanding how amorphicity influences electronic transport in nanowires, as they maintain translation invariance along their axis. As a consequence, the magnetic flux is constant along the nanowire and the surface recovers an effective time-reversal symmetry at $\phi=(2n+1)\phi_{0}/2$ that results in a perfectly transmitted mode.
@@
We demonstrate the presence of a perfectly transmitted mode by calculating the conductance of a single amorphicity realization as a function of the Fermi energy, shown in Fig. 1. In the absence of flux the conductance is zero until $E_{F}\simeq 0.06$, indicating the presence of an energy gap. At $\phi/\phi_{0}=0.56$ the conductance is quantized to $G=e^{2}/h$ for $E_{F}<0.12$, reflecting the perfectly transmitted mode. Note that the conductance is not exactly quantized down to $E_{F}=0$ because the leads have a finite size gap at $E_{F}=0$ when $\mu=0$. Since the surface state has a finite penetration length into the bulk, the flux it encloses deviates from the exact flux threaded through the full cross section. As a result, the perfectly transmitted mode does not occur exactly at $\phi/\phi_{0}=1/2$, but approaches this value in the limit of large $N_{x},N_{y}$  [^4] Apart from the shift induced by the finite bulk penetration depth, the position of the peaks may also be shifted to $\phi>(2n+1)\phi_{0}/2$ if the amorphous cross section is smaller than reference cross section area $\mathcal{A}_{\rm ref}$, and to $\phi<(2n+1)\phi_{0}/2$ if the amorphous cross section is bigger. This effect depends on the particular amorphicity realization.. Increasing the Fermi energy makes the flux value at which the conductance is maximal alternate between $\phi/\phi_{0}=0$ and $\phi/\phi_{0}=0.56$, as the number of modes at the Fermi energy changes between even and odd. The main difference from the crystalline case is the sharp conductance resonances seen especially at high Fermi energy. These come from the interface between the wire and the leads, where the lattice structure changes from being perfectly crystalline to being layered-amorphous, allowing electrons to backscatter and interfere. When the Fermi energy reaches the bands hosting states with a nonzero bulk contribution, scattering becomes stronger leading to an increase in interference and the upward trend in conductance gets flattened.
@@
In Fig. 2(a-c) we show the realization-averaged conductance of a layered-amorphous nanowire as a function of the Fermi energy and at three different amorphicities. As $w$ increases, the atomic arrangements of the different nanowire realizations become more distinct from each other.
@@
Figure 2: (a-c) Realization-averaged conductance of a layered-amorphous nanowire for three different amorphicities as a function of Fermi energy. The average is performed over 300300 layer amorphicity realizations of nanowires with $L=100$ and $N_{x}=N_{y}=10$. The chemical potential in the leads is $\mu=-1$. The shaded regions indicate the standard deviation of the distribution. (d) Single-realization conductance at $E_{F}=0$ as a function of magnetic flux and for different $w$. The parameters are $\mu=-1$, $L=200$ and $N_{x}=N_{y}=10$. (e) Band gap for the different nanowires in (d).
@@
Since the exact details of each conductance realization depend on the particular nanowire structure, conductance curves dephase when averaged, leading to smoother curves than for small $w$. The oscillations in the conductance at small $w$ are of the same nature as those at high Fermi energies in Fig. 1, although enhanced by the difference in chemical potential between leads and nanowire, and are present for each individual realization. Close to $E_{F}=0$, the perfectly transmitted mode dominates the conductance and is not affected by interference as it cannot backscatter. Fig. 2(d) shows the single-realization conductance at $E_{F}=0$ as a function of magnetic flux, and Fig. 2(e) the associated band gap. Conductance peaks coincide with gap closings, deviating from the crystalline case as either $w$ or $\phi$ increase. Despite these deviations, the conductance peaks reach perfect transmission up to $w=0.2$. Note that this is not the case if we calculate the average conductance since the flux at which perfect transmission is attained changes between realizations, decreasing the height of the peak.